Bug 1366749 - "Monitoring failed" messages during raid/mirror takeover
Summary: "Monitoring failed" messages during raid/mirror takeover
Keywords:
Status: CLOSED ERRATA
Alias: None
Product: Red Hat Enterprise Linux 7
Classification: Red Hat
Component: lvm2
Version: 7.3
Hardware: x86_64
OS: Linux
unspecified
medium
Target Milestone: rc
: ---
Assignee: LVM and device-mapper development team
QA Contact: cluster-qe@redhat.com
URL:
Whiteboard:
Depends On:
Blocks:
TreeView+ depends on / blocked
 
Reported: 2016-08-12 17:09 UTC by Corey Marthaler
Modified: 2016-11-04 04:17 UTC (History)
8 users (show)

Fixed In Version: lvm2-2.02.164-3.el7
Doc Type: No Doc Update
Doc Text:
Intra-release bug, no documentation needed.
Clone Of:
Environment:
Last Closed: 2016-11-04 04:17:27 UTC


Attachments (Terms of Use)


Links
System ID Priority Status Summary Last Updated
Red Hat Product Errata RHBA-2016:1445 normal SHIPPED_LIVE lvm2 bug fix and enhancement update 2016-11-03 13:46:41 UTC

Description Corey Marthaler 2016-08-12 17:09:23 UTC
Description of problem:

[root@host-082 ~]# lvs -a -o +devices
  LV                   VG            Attr       LSize   Pool Origin Data%  Meta%  Move Log Cpy%Sync Convert Devices                                    
  root                 rhel_host-082 -wi-ao----   6.67g                                                     /dev/vda2(205)                             
  swap                 rhel_host-082 -wi-ao---- 820.00m                                                     /dev/vda2(0)                               
  transform            test          rwi-a-r---  28.00m                                    100.00           transform_rimage_0(0),transform_rimage_1(0)
  [transform_rimage_0] test          iwi-aor---  28.00m                                                     /dev/sda1(1)                               
  [transform_rimage_1] test          iwi-aor---  28.00m                                                     /dev/sdb1(1)                               
  [transform_rmeta_0]  test          ewi-aor---   4.00m                                                     /dev/sda1(0)                               
  [transform_rmeta_1]  test          ewi-aor---   4.00m                                                     /dev/sdb1(0)

                               
[root@host-082 ~]# lvconvert --type mirror --mirrorlog disk test/transform
  Using default stripesize 64.00 KiB.
Are you sure you want to convert test/transform back to the older "mirror" type? [y/n]: y
  Monitoring test/transform failed.
  Logical volume test/transform successfully converted.


Is this a real "failure"? Or should this message be the expected "No longer monitoring ..." that you see normally when monitoring needs to be toggled off and on due to creates/removes/converts?



Version-Release number of selected component (if applicable):
3.10.0-489.el7.x86_64

lvm2-2.02.163-1.el7    BUILT: Wed Aug 10 06:53:21 CDT 2016
lvm2-libs-2.02.163-1.el7    BUILT: Wed Aug 10 06:53:21 CDT 2016
lvm2-cluster-2.02.163-1.el7    BUILT: Wed Aug 10 06:53:21 CDT 2016
device-mapper-1.02.133-1.el7    BUILT: Wed Aug 10 06:53:21 CDT 2016
device-mapper-libs-1.02.133-1.el7    BUILT: Wed Aug 10 06:53:21 CDT 2016
device-mapper-event-1.02.133-1.el7    BUILT: Wed Aug 10 06:53:21 CDT 2016
device-mapper-event-libs-1.02.133-1.el7    BUILT: Wed Aug 10 06:53:21 CDT 2016
device-mapper-persistent-data-0.6.3-1.el7    BUILT: Fri Jul 22 05:29:13 CDT 2016
cmirror-2.02.163-1.el7    BUILT: Wed Aug 10 06:53:21 CDT 2016
sanlock-3.4.0-1.el7    BUILT: Fri Jun 10 11:41:03 CDT 2016
sanlock-lib-3.4.0-1.el7    BUILT: Fri Jun 10 11:41:03 CDT 2016
lvm2-lockd-2.02.163-1.el7    BUILT: Wed Aug 10 06:53:21 CDT 2016

Comment 2 Alasdair Kergon 2016-08-22 17:18:50 UTC
Reproduced.

Comment 3 Alasdair Kergon 2016-08-23 17:46:03 UTC
Fixed upstream for next release.

When unmonitoring, the dso for the new segment type was supplied and failing to match the one for the old segment type, so nothing was changing, leaving the LV still monitored by the wrong dso at the end.

The code now finds out from dmeventd itself which dso is in use, and supplies that directly to the unmonitor call.

Comment 5 Corey Marthaler 2016-08-26 20:08:31 UTC
Fix verified in the latest rpms.

lvm2-2.02.164-3.el7    BUILT: Wed Aug 24 05:20:41 CDT 2016
lvm2-libs-2.02.164-3.el7    BUILT: Wed Aug 24 05:20:41 CDT 2016
lvm2-cluster-2.02.164-3.el7    BUILT: Wed Aug 24 05:20:41 CDT 2016


[root@host-117 ~]# lvcreate --type raid1 -L 50M -m 1 -n transform test
  Rounding up size to full physical extent 52.00 MiB
  Logical volume "transform" created.

[root@host-117 ~]# lvs -a -o +devices
  LV                   VG    Attr       LSize   Cpy%Sync Devices
  transform            test  rwi-a-r---  52.00m 100.00   transform_rimage_0(0),transform_rimage_1(0)
  [transform_rimage_0] test  iwi-aor---  52.00m          /dev/sda1(1)
  [transform_rimage_1] test  iwi-aor---  52.00m          /dev/sdb1(1)
  [transform_rmeta_0]  test  ewi-aor---   4.00m          /dev/sda1(0)
  [transform_rmeta_1]  test  ewi-aor---   4.00m          /dev/sdb1(0)

[root@host-117 ~]# lvconvert --type mirror test/transform
Are you sure you want to convert test/transform back to the older "mirror" type? [y/n]: y
  Logical volume test/transform successfully converted.

[root@host-117 ~]# lvs -a -o +devices
  LV                   VG   Attr       LSize   Log              Cpy%Sync Devices
  transform            test mwi-a-m---  52.00m [transform_mlog] 100.00   transform_mimage_0(0),transform_mimage_1(0)
  [transform_mimage_0] test iwi-aom---  52.00m                           /dev/sda1(1)
  [transform_mimage_1] test iwi-aom---  52.00m                           /dev/sdb1(1)
  [transform_mlog]     test lwn-aom---   4.00m                           /dev/sdg1(0)

[root@host-117 ~]# lvconvert --type raid1 test/transform
  Logical volume test/transform successfully converted.

[root@host-117 ~]# lvs -a -o +devices
  LV                   VG    Attr       LSize   Cpy%Sync Devices
  transform            test  rwi-a-r---  52.00m 100.00   transform_rimage_0(0),transform_rimage_1(0)
  [transform_rimage_0] test  iwi-aor---  52.00m          /dev/sda1(1)
  [transform_rimage_1] test  iwi-aor---  52.00m          /dev/sdb1(1)
  [transform_rmeta_0]  test  ewi-aor---   4.00m          /dev/sda1(0)
  [transform_rmeta_1]  test  ewi-aor---   4.00m          /dev/sdb1(0)

Comment 7 errata-xmlrpc 2016-11-04 04:17:27 UTC
Since the problem described in this bug report should be
resolved in a recent advisory, it has been closed with a
resolution of ERRATA.

For information on the advisory, and where to find the updated
files, follow the link below.

If the solution does not work for you, open a new bug report.

https://rhn.redhat.com/errata/RHBA-2016-1445.html


Note You need to log in before you can comment on or make changes to this bug.